# Principled priors for Bayesian inference of circular models

### v2

@@ 3 Methodology
This section introduces a principled framework for selecting priors for circular distributions. The framework aims at providing guidance so that the chosen priors could guard against producing over-complex models. The framework is given in Section 3.1 and structured around two essential steps: base model selection and distance computation. These two steps will be elaborated in the following subsections. On the basis of this framework, an illustrative procedure for deriving PC priors specifically designed for circular distributions will be presented afterwards.
3.1 Framework for Model Complexity Check
With no sufficient information supporting that the data require a complicated model, following the principle of Occam’s razor, it is necessary to examine the model complexity when setting a prior.
Statistical models usually have simpler counterparts, for example, having zero variance, approaching standard normal distribution, and so on. Therefore, it is reasonable to regard these special cases as the simplest models. Simpson et al. (2017) regards the simplest model of a corresponding distribution as the “base model”. In order not to increase the complexity of Bayesian model and to prevent overfitting issues, simpler Bayesian models are preferred when there is insufficient evidence supporting more complex models. One way to reduce the Bayesian model complexity is through selecting a proper prior distribution in order to make sure the prior has reasonable impact on the model.
One intuitive way to explain model complexity is to express it as “the distance between the model of interest and a corresponding base model”. We let $\pi\left(x\mid\xi\right)$ be the model of interest, where $x$ is the data and $\xi$ is the parameter of interest, and let $\pi\left(x\mid\xi_{0}\right)$ represent the base model, with $\xi=\xi_{0}$ be the corresponding parameter value at the base model. Simpson et al. (2017) proposed a reasonable distance measure, and in circular scenarios it can be written as
[…]
where $\text{KLD}\left(\pi\left(x\mid\xi\right)\big{\|}\pi\left(x\mid\xi_{0}\right)\right)$ is the Kullback-Leibler divergence (KLD) (Kullback & Leibler, 1951) between the model of interest and the base model.
Assume $\pi\left(\xi\right)$ is the prior for the parameter of interest, to examine whether a prior distribution prefers simpler models, we check the behavior of the prior density in distance scale given by
$\pi\left(d\right)$ $=\pi\left(\xi\left(d\right)\right)\left\lvert\frac{\partial\xi\left(d\right)}{\partial d}\right\rvert,$ (8)
where $\xi\left(d\right)$ is the inverse function of $d\left(\xi\right)$. If a prior $\pi\left(d\right)$ puts very low density at and around $d=0$, it indicates very low probability to obtain the simplest model. In other words, the prior prefers more complex models, which could lead to overfitting issues, especially when lacking sufficient information and belief on setting priors.
To prevent overfitting and reduce model complexity, priors that place a reasonable amount of density around $d=0$ and decay in the distance scale $\pi\left(d\right)$ as $d$ increases naturally become safer choices. However, to employ the proposed model complexity checking framework, the base model must be chosen properly. This is discussed in Section 3.2.
3.2 Base Models in Circular Scenarios
@@
In Euclidean space, statistical measures such as standard deviation and variance are intuitive and straightforward in terms of interpretation. However, these quantities lose their intuitive meanings for data collected on a circular domain (or transformed into circular form). To describe the spread or sparsity of circular data, most circular distributions introduce a “concentration” parameter, which quantifies the extent of concentration (or dispersion) within the data. The parameterization of this concentration parameter varies among circular distributions but typically includes the value “0” as a special case, representing no concentration (or maximum dispersion). In most common cases, the distribution is simplified to the circular uniform distribution.
@@
To conclude, the two boundaries of the support of the concentration parameters of circular distributions typically correspond to two distinct special cases: one representing maximum dispersion (e.g., the circular uniform distribution) and the other representing maximum concentration (e.g., a point mass). These boundary cases are inherently simpler than intermediate cases, making them natural candidates for base models.
In addition, it is worth noticing that there is no natural choice of base model for the location parameter $\mu$ of circular distributions as it represents the mean direction of the data on a unit circular ($\mu\in\left[0,2\pi\right)$). Therefore, the models with different $\mu$ values share the same model complexity. Based on the nature of mean parameter of circular distributions, it is reasonable and recommended to set circular uniform prior in advance since the model complexity is equal everywhere.
3.3 General Procedure for deriving PC Prior
Examining the priors’ behavior in distance scale with the framework given in Section 3.1 is important in order to prevent overfitting. However, this process is tedious and inefficient since the base model, the distance and the prior density in distance scale differ for different distributions and priors. To simplify the prior selection process and to prevent inconsistency of the prior performance caused by the choice of base models and priors, we suggest to employ the PC prior for circular Bayesian analysis. We propose the general procedure for deriving PC priors for circular distributions in this section, together with the exact expression of PC priors for the most popular circular distributions in Section 4.
@@ 3.3 General Procedure for deriving PC Prior
Following the formulation of Simpson et al. (2017), the Penalized Complexity (PC) prior for a model parameter $\xi$ is defined by a probability density function as follows:
@@
$\pi\left(d\right)=\lambda\exp\left\{-\lambda d\right\}\Longleftrightarrow\pi\left(\xi\right)=\lambda\exp\left\{-\lambda d\left(\xi\right)\right\}\left\lvert\frac{\partial d\left(\xi\right)}{\partial\xi}\right\rvert.$ (9)
where $\lambda$ is the scaling parameter of the PC prior. The PC prior inherently favors simpler models, penalizing prior density exponentially as the distance $d\left(\xi\right)$ increases.
The choice of the base model and the scaling parameter $\lambda$ are critical both to the derivation and the performance of PC priors. These aspects are explored in Section 3.2 and Section 3.4, respectively.
3.4 Strategies for Setting Scaling Parameter
@@
The scaling parameter, $\lambda$, of the PC prior determines the strength of penalization for deviations away from the base model. To appropriately select the scaling parameter, users need to have some understanding of the data and its properties.
Simpson et al. (2017) proposed selecting the scaling parameter by:
$P_{\mathcal{PC}}\left(Q\left(\xi\right)>U\right)$ $=\alpha,$ (10)
where $U$ represents the threshold for a “tail event” as perceived by the user, and $\alpha$ is the probability assigned to this event. This approach involves transforming the parameter of interest, $\xi$, into an interpretable form $Q\left(\xi\right)$, allowing users to prescribe the informativeness of the PC prior by controlling the weight assigned to tail events.
@@
For most circular distributions, where the bounds of the concentration parameter $\xi$ correspond to the circular uniform distribution and a point mass, a suitable transformation $Q\left(\xi\right)$ can often be found. For instance, one could define $Q\left(\xi\right)$ such that circular uniform density corresponds to $Q\left(\xi\right)\to 2\pi$, and the point mass corresponds to $Q\left(\xi\right)\to 0$. Under this transformation, $Q\left(\xi\right)\in\left(0,2\pi\right)$, making Equation 10 represent “the probability that the data fall outside a radian $U$ around the mean direction being equal to $\alpha$. The small values of $U$ indicate a high concentration to a point mass, whilst large values of $U$ mean that the user believes the data spread widely. This transformation is intuitive and allows users to gain insights into appropriate values for $U$ and $\alpha$ by visualizing the data.
4 PC Priors for Popular Circular Distributions
Section 3 outlines the general framework for selecting proper priors tailored to circular distributions and the procedure for deriving the associated PC priors. In this section, we present PC priors for some of the widely used circular distributions discussed in Section 2. In addition, based on the discussion in Section 3.2, instead of the position parameter (mean $\mu$), we focus on the PC priors for the concentration parameter.
4.1 PC Prior for Von Mises Distribution
@@
The first step in deriving the PC prior is to compute the KLD (Section 3.3). For the vM distribution as defined in Equation 2, the KLD is given below:
[…]
@@
Recall the behavior of the vM density with respect to changes in the concentration parameter $\kappa$ (Section 2.1). Two natural choices for base model for the concentration parameter are the circular uniform distribution ($\kappa_{0}=0$), representing no concentration, and the point mass ($\kappa_{0}\gg\kappa$), representing high concentration. With these two chosen base models, the corresponding PC priors for $\kappa$ can be computed and are given in Proposition 4.1 and Proposition 4.2 respectively. The proofs are given in Section A.1.
@@
$\pi(\kappa)$ $=\lambda\exp\left\{-\lambda\sqrt{\frac{\kappa\mathcal{I}_{1}(\kappa)}{\mathcal{I}_{0}(\kappa)}-\log\left(\mathcal{I}_{0}\left(\kappa\right)\right)}\right\}\frac{\frac{\kappa\left(\mathcal{I}_{0}\left(\kappa\right)+\mathcal{I}_{2}\left(\kappa\right)\right)}{2\mathcal{I}_{0}\left(\kappa\right)}-\frac{\kappa\mathcal{I}_{1}\left(\kappa\right)^{2}}{\mathcal{I}_{0}\left(\kappa\right)^{2}}}{2\sqrt{\frac{\kappa\mathcal{I}_{1}(\kappa)}{\mathcal{I}_{0}(\kappa)}-\log\left(\mathcal{I}_{0}\left(\kappa\right)\right)}},$ (12)
@@
$\pi\left(\kappa\right)$ $=\lambda\exp\left\{-\lambda\sqrt{1-\frac{\mathcal{I}_{1}(\kappa)}{\mathcal{I}_{0}(\kappa)}}\right\}\frac{\left(\frac{\mathcal{I}_{0}(\kappa)+\mathcal{I}_{2}(\kappa)}{2\mathcal{I}_{0}(\kappa)}-\frac{\mathcal{I}_{1}(\kappa)^{2}}{\mathcal{I}_{0}(\kappa)^{2}}\right)}{2\sqrt{1-\frac{\mathcal{I}_{1}(\kappa)}{\mathcal{I}_{0}(\kappa)}}},$ (14)
@@
Figure 4: $\kappa_{0}=0$ base model PC prior density ($\pi\left(\kappa\right)$, left) and log-log density ($\log\pi\left(\log\kappa\right)$, right).
Figure 5: $\kappa_{0}\rightarrow\infty$ base model PC prior density ($\pi\left(\kappa\right)$, left) and log-log density ($\log\pi\left(\log\kappa\right)$, right).
4.1.1 Interpretable Parameter Transformation
@@
$\lambda=\begin{cases}-\frac{\log\left(1-\alpha\right)}{\sqrt{\frac{\left(\frac{2\pi}{U}-1\right)\mathcal{I}_{1}(\frac{2\pi}{U}-1)}{\mathcal{I}_{0}(\frac{2\pi}{U}-1)}-\log\left(\mathcal{I}_{0}\left(\frac{2\pi}{U}-1\right)\right)}}\quad&\text{for base model at}\quad\kappa_{0}=0;\\ -\frac{\log\left(1-\alpha\right)}{\sqrt{1-\frac{\mathcal{I}_{1}(\frac{2\pi}{U}-1)}{\mathcal{I}_{0}(\frac{2\pi}{U}-1)}}}\quad&\text{for base model at}\quad\kappa_{0}\rightarrow\infty.\end{cases}$ (17)
4.2 PC Prior for Cardioid Distribution
For the cardioid distribution as defined in Equation 5, we have the KLD as
@@ 4.2 PC Prior for Cardioid Distribution
The base models for the concentration parameter $\ell$ of cardioid distribution are unusual, as $\ell$ controls the extent of the cardioid curve deviating away from the circular uniform distribution. Therefore, the two base models for $\ell$ are the circular uniform distribution $\ell_{0}=0$ and the most un-uniform cardioid curve $\left(\ell_{0}\rightarrow\frac{1}{2}\right)$, and the corresponding PC prior for $\ell$ are given in Proposition 4.3 and Proposition 4.4. The proofs are given in Section A.2.
@@
$\pi\left(\ell\right)$ $=\lambda\exp\left\{-\lambda d\left(\ell\right)\right\}\frac{2\ell}{\left(1-\exp\left\{-\lambda\sqrt{1-\log\left(2\right)}\right\}\right)\left(1+\sqrt{1-4\ell^{2}}\right)d\left(\ell\right)},$ (19)
@@
$\pi\left(\ell\right)$ $=\lambda\exp\left\{-\lambda d\left(\ell\right)\right\}\frac{2\ell+\sqrt{1-4\ell^{2}}-1}{2\ell d\left(\ell\right)},$ (22)
@@
Figure 6: $\ell_{0}=0$ base model PC prior density ($\pi\left(\ell\right)$, left) and log-log density ($\log\pi\left(\log\ell\right)$, right).
Figure 7: $\ell_{0}\rightarrow 0.5$ base model PC prior density ($\pi\left(\ell\right)$, left) and log-log density ($\log\pi\left(\log\ell\right)$, right).
4.2.1 Interpretable Parameter Transformation
@@
$\lambda=\begin{cases}-\frac{\log\left(\alpha+\left(1-\alpha\right)\exp\left\{-\lambda\sqrt{1-\log 2}\right\}\right)}{d\left(\frac{U}{2}\right)}\quad&\text{for base model at}\quad\ell_{0}=0;\\ -\frac{\log\left(1-\alpha\right)}{d\left(\frac{U}{2}\right)}\quad&\text{for base model at}\quad\ell_{0}\rightarrow\frac{1}{2},\end{cases}$ (26)
@@
4.3 PC Prior for Wrapped Cauchy Distribution
Recall that $\rho=1$ is not well defined for the wrapped Cauchy density (Section 2.3), and also that the WC distribution is preferred as a model with less concentration, when compared with vM distribution. WC distribution is thus usually employed when we believe the data are not strongly concentrated. Thus, it is more reasonable to consider the circular uniform distribution ($\rho_{0}=0$) as the only base model for concentration parameter $\rho$ of the WC distribution. Therefore, the KLD between $\pi\left(x\mid\mu,\rho\right)$ and $\pi\left(x\mid\mu,\rho_{0}=0\right)$ is
[…]

### v3

@@ 3 Methodology
This section introduces a principled framework for constructing hyperpriors for circular distributions that favors simpler circular models. The guiding principles, along with the corresponding prior formulation framework that satisfies them, are presented in Section 3.1, and specific priors for widely used circular distributions are proposed in Section 3.2.
3.1 Penalizing complexity prior for circular models
When there is no sufficient information supporting that the data require a complicated model, following the principle of Occam’s razor, it is necessary to examine model complexity when setting a prior. One effective way to penalize Bayesian model complexity is through the selection of a proper prior distribution, ensuring that the prior exerts a reasonable influence on the model’s complexity. With this motivation, Simpson et al. (2017) proposed the PC prior framework, built upon four principles: Occam’s razor, measure of complexity, constant rate penalization and user-defined scaling. The measure of complexity principle states that the prior should be constructed using a model complexity measure defined by the “distance” $d$ between the model of interest and a simpler base model; The constant rate penalization principle specifies that the PC prior density decays at a constant rate $r\in\left(0,1\right)$ with respect to an increment $\delta$ in the distance $d$, satisfying
$\frac{p_{d}\left(d+\delta\right)}{p_{d}\left(d\right)}=r^{\delta},\quad d,\delta\geq 0.$ (7)
Finally, the user-defined scaling principle suggests that users should have at least some understanding of the data, or of how complex the model is expected to be, to define the strength of penalization accordingly.
The PC prior framework has been successfully applied across a wide range of fields, demonstrating both its practicality and robustness. For instance, it has been widely used in spatial and spatio-temporal modelling (Fuglstad et al., 2019; Cabral et al., 2023; Rodriguez Avellaneda et al., 2025), time-series analysis (Sørbye and Rue, 2017), survival modelling (Van Niekerk et al., 2021), and nonparametric regression (Ventrucci and Rue, 2016). These applications highlight the flexibility of PC priors in controlling model complexity across diverse modelling contexts, motivating their extension to circular data.
@@ 3.1 Penalizing complexity prior for circular models
In Euclidean space, statistical measures such as standard deviation and variance are intuitive and straightforward in terms of interpretability. However, as discussed in Section 2, these quantities lose their intuitive meanings for data collected on a circular domain (or transformed into circular form) and concentration parameters are used. The parameterization of the concentration parameter varies among circular distributions but typically includes the value “0” as a special case, representing no concentration (or maximum dispersion). In most common cases, the distribution is simplified to the circular uniform distribution. It should be noted that the mean direction is not likelihood identifiable in the circular uniform distribution.
@@
Although most circular distributions exhibit desirable properties when their concentration parameters approach the boundary values, the inference for these distributions can be unstable because the data are defined on a limited range ($0,2\pi)$). As a result, small variations in the data can correspond to large changes on the circular scale, making the concentration parameter highly sensitive. In particular, when the data deviate from the boundary cases, that is, when they are neither strongly clustered nor close to circular uniformity, the underlying model structure becomes more complex and the estimation of the concentration parameters becomes difficult and less stable.
This sensitivity and instability highlight the need for a principled approach to regularizing model complexity. One natural way to achieve this is to view the boundary cases as simpler circular models, which can be regarded as the “base model” in the PC prior framework, that is, the simplest models representing minimal structure and the most straightforward interpretation. The notion of model complexity can then be described by the measure of complexity principle, where the model complexity is quantified as the distance from the base model. We let $p\left(x\mid\xi\right)$ be the model of interest, where $x$ is the data and $\xi$ is the parameter of interest, and let $p\left(x\mid\xi_{0}\right)$ represent the base model, with $\xi=\xi_{0}$ be the corresponding parameter value at the base model. Simpson et al. (2017) proposed a reasonable distance measure, and in circular scenarios it can be written as
$d\left(\xi\right)$ $=\sqrt{\text{KLD}\left(p\left(x\mid\xi\right)\big\|p\left(x\mid\xi_{0}\right)\right)}=\sqrt{\int_{0}^{2\pi}p\left(x\mid\xi\right)\log\left(\frac{p\left(x\mid\xi\right)}{p\left(x\mid\xi_{0}\right)}\right)dx},$ (8)
where $\text{KLD}\left(\cdot\big\|\cdot\right)$ denotes the Kullback-Leibler divergence (KLD) (Kullback and Leibler, 1951) between two distributions, which is one of the most widely used quantitative measures of the difference between probability distributions.
@@
Following the formulation of Simpson et al. (2017), the Penalized Complexity (PC) prior for a model parameter $\xi$ is defined as an exponential probability density function with respect to the distance, as follows:
@@
$p\left(d\right)=\lambda\exp\left\{-\lambda d\right\}\Longleftrightarrow p\left(\xi\right)=\lambda\exp\left\{-\lambda d\left(\xi\right)\right\}\left\lvert\frac{\partial d\left(\xi\right)}{\partial\xi}\right\rvert$. (9)
where $\lambda$ is the scaling parameter of the PC prior, defined such that $P_{\mathcal{PC}}\left(Q\left(\xi\right)>U\right)=\alpha$, with $Q\left(\xi\right)$ being a function of $\xi$, $U$ being a reasonable value of $Q\left(\xi\right)$ and $\alpha\in\left(0,1\right)$ a value representing probability. The formulations of $Q\left(\xi\right)$ are discussed later in this section.
To illustrate how the PC prior behaves with respect to the distance measure, Figure 4 is presented. The left-hand-side plot shows the distance $d\left(\kappa\right)$ between the von Mises distribution and the circular uniform distribution. If we express the prior density for $\kappa$, $p\left(\kappa\right)$, in terms of distance, $p\left(d\right)$, then from the right-hand-side plot of Figure 4, we can observe that the prior density decreases as the distance increases. Since PC priors are exponential with respect to distance, for any concentration parameter $\xi$ of a circular distribution, the density of its PC prior will always be penalized as the distance between that circular distribution and the base model increases. Moreover, because the prior is defined to be exponential on the distance scale, this penalization occurs at a constant rate with respect to increments in distance. Therefore, the PC prior satisfies the constant rate penalization principle, and inherently favors simpler models, penalizing the prior density exponentially as the distance $d(\xi)$ increases.
Figure 4: Distance ($d\left(\kappa\right)$) between vM distribution and circular uniform distribution (left) and the PC prior density in distance scale for different $\lambda$ values (right).
@@
The scaling parameter, $\lambda$, of the PC prior determines the strength of penalization for deviations away from the base model. To satisfy the ser-defined scaling principle and to appropriately select the scaling parameter, users need to have some understanding of the data and its properties. A general strategy for selecting $\lambda$ is to set $Q\left(\xi\right)=\psi$ , where $\psi\in\left[0,1\right]$ denotes the mean resultant length. Following Ley and Verdebout (2017) and Lund et al. (2017), it is defined as
$\psi$ $=\frac{1}{n}\sqrt{\left(\sum_{i=1}^{n}\cos\left(x_{i}\right)\right)^{2}+\left(\sum_{i=1}^{n}\sin\left(x_{i}\right)\right)^{2}},$ (10)
for data $x_{i},i=1,2,\cdots,n$. Alternative definitions of this quantity also exist, such as moving the factor $\frac{1}{n}$ inside the two squared summations (Mardia and Jupp, 2009). This quantity is a measure of the concentration around the circle. A value of $\psi=0$ corresponds to a circular uniform distribution where there is no preferred direction, while $\psi=1$ indicates that all observations coincide at the same angle (maximum concentration). Under this interpretation, $P_{\mathcal{PC}}\left(Q\left(\xi\right)>U\right)=\alpha$ expresses the belief that the probability that the mean resultant length of the data exceeds $U$, is $\alpha$.
This interpretation is model-independent, as the user can also express $\psi$ explicitly as a function of the model parameter $\xi$, and therefore interpret $Q\left(\xi\right)$ in terms of $\xi$. For instance, in the von Mises distribution, $\psi\left(\kappa\right)=\frac{I_{1}(\kappa)}{I_{0}(\kappa)}$, whilst for the cardioid and wrapped Cauchy distributions, $\psi$ equals to the concentration parameters $\ell$ and $\rho$ respectively. Conceptually, the mean resultant length serves as a sample-based indicator of data concentration, while the concentration parameter of a circular distribution controls the shrinkage behavior of the corresponding probability density. Therefore, the mean resultant length links the circular distribution to the observed data, incorporating the information from the data into the density function.
This unified approach covers various circular models under one framework. However, it may not be intuitive for users inexperienced with circular data. Therefore, an alternative, more intuitive approach is also suggested.
@@
For most circular distributions, where the bounds of the concentration parameter $\xi$ correspond to the circular uniform distribution and a point mass, a suitable transformation $Q\left(\xi\right)$ can often be found. For instance, one could define $Q\left(\xi\right)$ such that circular uniform density corresponds to $Q\left(\xi\right)\to 2\pi$, and the point mass corresponds to $Q\left(\xi\right)\to 0$. Under this transformation, $Q\left(\xi\right)\in\left(0,2\pi\right)$, making $P_{\mathcal{PC}}\left(Q\left(\xi\right)>U\right)=\alpha$ represent “the probability that the data fall outside a radian $U$ around the mean direction being equal to $\alpha$. The small values of $U$ indicate a high concentration to a point mass, whilst large values of $U$ mean that the user believes the data spread widely. This transformation is intuitive and allows users to gain insights into appropriate values for $U$ and $\alpha$ by visualizing the data.
It is worth noting that the proposed PC priors are flexible in terms of informativeness. That is, they can be either weakly or strongly informative (Simpson et al., 2017), depending on the user’s belief about the data. As an illustration, consider selecting the value of $\lambda$ using the mean resultant length ($Q(\xi)=\psi$). If the user believes that “the mean resultant length of the data is nearly impossible to be higher than 0.8,” then one can set $U=0.8$ and $\alpha=0.001$ to obtain a strongly informative prior. Conversely, if the user has no prior knowledge about the concentration of the data, $\lambda$ can be chosen by setting $U=0.5$ and $\alpha=0.5$, indicating uncertainty about whether the mean resultant length of the data should be larger or smaller than the median, which consequently yields a weakly informative prior. Therefore, the PC prior can always be employed regardless of whether sufficient prior information is available.
3.2 Specific cases
3.2.1 Von Mises Distribution
@@
The first step in deriving the PC prior is to compute the KLD. For the vM distribution as defined in Equation 1, the KLD is given below:
$\text{KLD}\left(\mathcal{VM}\big\|\mathcal{VM}_{0}\right)$ $=\int_{0}^{2\pi}p\left(x\mid\mu,\kappa\right)\log\left(\frac{p\left(x\mid\mu,\kappa\right)}{p\left(x\mid\mu,\kappa_{0}\right)}\right)dx$ (11)
@@
Recall the behavior of the vM density with respect to changes in the concentration parameter $\kappa$ (Section 2.1). Two natural choices for base model for the concentration parameter are the circular uniform distribution ($\kappa_{0}=0$), representing no concentration, and the point mass ($\kappa_{0}\gg\kappa$), representing high concentration. With these two chosen base models, the corresponding PC priors for $\kappa$ can be computed and are given in Proposition 3.1 and Proposition 3.2 respectively. The proofs are given in Appendix A.1.
@@
$p(\kappa)$ $=\lambda\exp\left\{-\lambda\sqrt{\frac{\kappa\mathcal{I}_{1}(\kappa)}{\mathcal{I}_{0}(\kappa)}-\log\left(\mathcal{I}_{0}\left(\kappa\right)\right)}\right\}\frac{\frac{\kappa\left(\mathcal{I}_{0}\left(\kappa\right)+\mathcal{I}_{2}\left(\kappa\right)\right)}{2\mathcal{I}_{0}\left(\kappa\right)}-\frac{\kappa\mathcal{I}_{1}\left(\kappa\right)^{2}}{\mathcal{I}_{0}\left(\kappa\right)^{2}}}{2\sqrt{\frac{\kappa\mathcal{I}_{1}(\kappa)}{\mathcal{I}_{0}(\kappa)}-\log\left(\mathcal{I}_{0}\left(\kappa\right)\right)}},$ (12)
@@
$p\left(\kappa\right)$ $=\lambda\exp\left\{-\lambda\sqrt{1-\frac{\mathcal{I}_{1}(\kappa)}{\mathcal{I}_{0}(\kappa)}}\right\}\frac{\frac{\mathcal{I}_{0}(\kappa)+\mathcal{I}_{2}(\kappa)}{2\mathcal{I}_{0}(\kappa)}-\frac{\mathcal{I}_{1}(\kappa)^{2}}{\mathcal{I}_{0}(\kappa)^{2}}}{2\sqrt{1-\frac{\mathcal{I}_{1}(\kappa)}{\mathcal{I}_{0}(\kappa)}}},$ (14)
@@
Figure 5: $\kappa_{0}=0$ base model PC prior density (left) and log-log density (right).
Figure 6: $\kappa_{0}\rightarrow\infty$ base model PC prior density (left) and log-log density (right).
@@
3.2.2 Cardioid Distribution
For the cardioid distribution as defined in (5), we have the KLD as
$\text{KLD}\left(\mathcal{C}\big\|\mathcal{C}_{0}\right)$ $=\int_{0}^{2\pi}p\left(x\mid\mu,\ell\right)\log\left(\frac{p\left(x\mid\mu,\ell\right)}{p\left(x\mid\mu,\ell_{0}\right)}\right)dx$ (18)
@@ 3.2.2 Cardioid Distribution
The base models for the concentration parameter $\ell$ of cardioid distribution are unusual, as $\ell$ controls the extent of the cardioid curve deviating away from the circular uniform distribution. Therefore, the two base models for $\ell$ are the circular uniform distribution $\ell_{0}=0$ and the most un-uniform cardioid curve $\left(\ell_{0}\rightarrow\frac{1}{2}\right)$, and the corresponding PC prior for $\ell$ are given in Proposition 3.3 and Proposition 3.4. The proofs are given in Appendix A.2.
@@
$p\left(\ell\right)$ $=\lambda\exp\left\{-\lambda d\left(\ell\right)\right\}\frac{2\ell}{\left(1-\exp\left\{-\lambda\sqrt{1-\log\left(2\right)}\right\}\right)\left(1+\sqrt{1-4\ell^{2}}\right)d\left(\ell\right)},$ (19)
@@
$p\left(\ell\right)$ $=\lambda\exp\left\{-\lambda d\left(\ell\right)\right\}\frac{2\ell+\sqrt{1-4\ell^{2}}-1}{2\ell d\left(\ell\right)},$ (22)
@@
Figure 7: $\ell_{0}=0$ base model PC prior density (left) and log-log density (right).
Figure 8: $\ell_{0}\rightarrow 0.5$ base model PC prior density (left) and log-log density (right).
@@
[…]
@@
3.2.3 Wrapped Cauchy Distribution
Recall that $\rho=1$ is not well defined for the wrapped Cauchy density (Section 2.3), and also that the WC distribution is preferred as a model with less concentration, when compared with vM distribution. WC distribution is thus usually employed when we believe the data are not strongly concentrated. Thus, it is more reasonable to consider the circular uniform distribution ($\rho_{0}=0$) as the only base model for concentration parameter $\rho$ of the WC distribution. Therefore, the KLD between $p\left(x\mid\mu,\rho\right)$ and $p\left(x\mid\mu,\rho_{0}=0\right)$ is
$\text{KLD}\left(\mathcal{WC}\big\|\mathcal{WC}_{0}\right)$ $=\int p\left(x\mid\mu,\rho\right)\log\left(\frac{p\left(x\mid\mu,\rho\right)}{p\left(x\mid\mu,\rho_{0}\right)}\right)dx$ (27)